OpenStudy (anonymous):

Determine all pairs of polynomials f and g with real coefficients such that \(x^2g(x)=f(g(x))\)

OpenStudy (kc_kennylau):

Let g(x) be a0+a1x+a2x^2+a3x^3+...+anx^n. Let f(x) be b0+b1x+b2x^2+b3x^3+...+bnx^n. a0x^2+a1x^3+a2x^4+...+anx^(n+2) = f(a0+a1x+a2x^2+...+anx^n) a0x^2+a1x^3+a2x^4+...+anx^(n+2) = b0 + b1(a0+a1x+a2x^2+...+anx^n) + b2(a0+a1x+a2x^2+...+anx^n)^2 + b3(a0+a1x+a2x^2+...+anx^n)^3 + ... + bn(a0+a1x+a2x^2+...+anx^n)^n Nah just kidding, wrong approach here.

OpenStudy (anonymous):

ok ;)

OpenStudy (kc_kennylau):

Maybe using calculus? x^2g(x) = f(g(x)) 2xg(x)+x^2g'(x) = f'(g(x))g'(x) 2xg(x) = f'(g(x))g'(x) - x^2g'(x) 2xg(x) = g'(x)[f'(g(x))-x^2] Nah this is not going the right way...

OpenStudy (kc_kennylau):

How do I separate f and g?! :O

OpenStudy (anonymous):

jus thinking about making a new function \(h(x)=(fog)(x)=f(g(x))\)

OpenStudy (kc_kennylau):

What's the difference...

OpenStudy (kc_kennylau):

f(g(x)) = h(x) g(x) = f^-1(h(x)) mother of G-d...

OpenStudy (kc_kennylau):

x^2f^-1(h(x))=h(x)

OpenStudy (kc_kennylau):

Still a chain function here...

OpenStudy (kc_kennylau):

@ganeshie8

ganeshie8 (ganeshie8):

no clue... still thinking..

ganeshie8 (ganeshie8):

first pair : g = x f = x^3 seem to satisfy

OpenStudy (anonymous):

maybe \(f(x)=x^{n-2},g(x)=x^n\)

OpenStudy (kc_kennylau):

Let f's order be m and g's order be n. O(x^2g(x)) = O(x^2)+O(g(x)) = n+2 O(f(g(x)) = O(f(x))O(g(x)) = mn x^2g(x)=f(g(x)) O[x^2g(x)]=O[f(g(x))] n+2=mn m=(n+2)/n n+2=mn mn-n=2 n(m-1)=2 n=2/(m-1)

OpenStudy (kc_kennylau):

Note that m and n must be integers.

ganeshie8 (ganeshie8):

looks great ! so m = 2, 3

OpenStudy (kc_kennylau):

Thanks @ganeshie8 for making me come to the answer.

OpenStudy (kc_kennylau):

And @Jonask too.

OpenStudy (kc_kennylau):

After some trial-and-errors we can say that n can be 1 or 2

OpenStudy (anonymous):

ok looks very good so far,the order's are working out properly

OpenStudy (kc_kennylau):

And as ganeshie said m can be 2 or 3

OpenStudy (kc_kennylau):

A million thanks to @ganeshie8 and a billion of them to @Jonask for asking this question :)

OpenStudy (kc_kennylau):

When m=2, n=2; When m=3, n=1.

ganeshie8 (ganeshie8):

how to conclude # of pairs of functions... there can be many functions wid those degree... or im not seeing something obvious uhmmm

OpenStudy (kc_kennylau):

When m=2 and n=2: Let f(x) and g(x) be ax^2+bx+c and dx^2+ex+f respectively. x^2g(x)=f(g(x)) x^2(dx^2+ex+f)=a(dx^2+ex+f)^2+b(dx^2+ex+f)+c blah blah blah

OpenStudy (anonymous):

m,n has to be integers ,so that makes only two pairs

OpenStudy (kc_kennylau):

When m=3 and n=1: Let f(x) and g(x) be ax^3+bx^2+cx+d and ex+f respectively. x^2g(x)=f(g(x)) x^2(ex+f)=a(ex+f)^3+b(ex+f)^2+c(ex+f)+d blah blah blah

OpenStudy (kc_kennylau):

These two pairs are actually solvable, but I'm looking for an easier way to find out all the pairs.

OpenStudy (anonymous):

a trillion thanks to @kc_kennylau for coming up with the "orders"

OpenStudy (anonymous):

6 variables thats not gonna be smooth

OpenStudy (anonymous):

but this could help \(0=f(g(0)),g(1)=f(g(1)\dots\)

OpenStudy (kc_kennylau):

And g(-1)=f(g(-1))

OpenStudy (anonymous):

so basically we can get all variables with this conditions,i am looking for a computerised solution

OpenStudy (kc_kennylau):

@ganeshie8 any idea?